Daily staff Journal or Duty Officers' Log

3d Sqdn, 4th Cav

XT654145 - 0001 31 Jan 68 to 2400 31 Jan 68

0035 SC15 spotted movement in the wood line at Hoc Mon. Fired a warning shot and 1 old man wounded.
0045 SA and SB SITREP No Change.
0150 SA, SB, and SC15 SITREP No Change.
0253 Fm Pewter Brigade 8. In cont w/EST 2 Co VC. Request LFT and flareship at XT359253. Cont division. They will handle.
0255 SB SITREP No Change.. SC10 SITREP No Change.
0307 Sounded mortar attack siren at 0306H.
0314 SA, SB, SC, SH, and LRRP - all personnel accounted for, neg casualties at this time.
0324 Call DTOC - DTOC Rear - Discom notified of status of Saber elements.
0327 Centaur feels they took many rounds in their area of aircraft.
0335 SB20 and SB30 LP's back in AA.
0337 SB elements in field - SITREP No Change.
0340 SB65Z. Reports all personnel accounted for - Negative hits in B area.
0341 SA - no hits in area.
0342 SC - Undetermined number of rounds in area.
0350 SA, SB, SC, Centaur, SH, and LRRP notified of stand-down. Are now mustering troops for count.
0400 SA SITREP No Change.
0401 SB reports no casualties; often no troop damage (muster).
0403 Centaur reports no casualties after muster

End of page 1

0405 SA reports no casualties after Muster; no damage.
0406 SC reports no casualties after muster, no damage.
0407 Hq and LRRP, negative casualties after muster.
0409 SC had 5 rounds hit in SC area - no damage.
0415 SC notified of poss reaction force to Hoc-Mon to block VC Regiment escape, to results ar attack on Ton-Son-Nhut AFB. VC may be dressed as ARVNs.
0416 Centaur has 1 minor casualty, not caused by shrapnel, 1 cut in foot, running to bunker.
0417 Lt Kelly is sending a piece of shrapnel that came from a rocket to the TOC.
0425 SC notified to mount vehicles and standby until further notice. SC6 notified. They will come to the TOC for briefing.
0432 BC under 2nd mortar attack.
0440 SC personnel on track. Centaur personnel are present and accounted for.  SA personnel accounted for. LRRP personnel accounted for. SH personnel accounted for.
0445 SC15 and SB SITREP No Change.  Warrior notified that SC will be moving out to escort a blocking force.
0450 Flareship is to be airborne 05.
0445 Trang-Bang is under attack.  SB10 receiving sniper fire. DTOC notified.
0457 Stand-Down 0456H, SA, SB, Sc, Cen, SH, LRRP.
0501 Flareship is airborne and orbiting awaiting for Saber6.

End of page 2

0510  SH personnel accounted for.
0511 Sa Stand-to complete.
0512 SC6 and elements on move at this time.
0517 SC elements cleared the Main Gate3. SC status is 3 Officers, 113 EM, 17 PCs, 4 tanks, and 2 M106's. SA status is 4 Officers, 106 EM, 7 tanks, 18 PCs, 1 M106, and 1 VTR.
0532 Mortar alart sounded 3rd time.
0534 SB stand-to complete.
0541 SH, SA elements accounted for.
0602 SC15 SITREP No Change.
0605 SA, SB sweep to Cu-Chi and C.P. 31.
0608 Stand-down, all clear.
0610 SC C.P. 'D'.
0615 EOD notified reference 2 duds at Delta in coral.
0618 SC lead C.P. 'E', trail C.P. 'D'.
0620 SA, SB holding sweeps.
0625 SC C.P. 'L'
0633 SC30 cleared C.P. 'E' to catch up with the rest of Saber elements.
0635 Gave DISCOM appraisal of damage.

End of page 3

0645 SC30 C.P. 'L'; closed with remainder of SC elements
0645 SC to go to XT795943 when unknown VC are at EST location. They have rockets
0650 SC15 moving from Hoc-Mon to join troop.
0653 SA10 is beginning sweep from C.P. 178 to TBB. SB20 to sweep and BB to Cu-Chi. B sweep elements will hold at horse Shoe until further notice.
0700 SC is approaching Ton-Son-Nhut; will go to westside and support SC working on new push, 58.25
0702 SC is approaching gate and TSN
0705 SB20 departed AA for sweep.
0706 SA10 departed C.P. 178 for sweep.
0708 SC is in contact this time and are taking heavy B40 rounds.
0710 SC6 has been hit. We are scrambling a gunteam at this time.
0714 SC elements from Hoc-Mon have brought up to SC elements.
0716 SA sweep at C.P. 2.
0718 SC15 entering the Main Gate
0720 SC20 at approximately XS740980
0723 1 dustoff airborne to SC location.
0725 A 1/8th will be ours again today.
0725 SB25 roadblocks at XT540178 at o715H, 2 RPG between dirt mounds and bushes; will clear tomorrow.

End of page 4

0724 SA10 C.P. 5.
0730 SC31 app 7896
0740 S6 ship went down. Centaur14
0742 SA and SB to bring troops to high state of readiness; also, SA sweep will remain at C.P. 31
0745 SC along MSR 1 still in heavy contact with and VC are trying to outflank from SW
0745 SB given a mission to move to TSN; also, SA is to deploy the other element not needed
0755 SB20 C.P. 21. B30 and 10 C.P.8.
0756 SA departing C.P. 178.
0800 S6 ship back and airborne this time; ship took hit, but not bad.
0800 Vindicator69 says artillery will contact us.
0803 SA C.P.2.
0805 S6 reports friendly ground forces coming up from the rear.  SC does not wish additional help this time.
0806 SB20 by-passing road block. (Bao-Trai).
0807 SB elements to hold C.P. 22.
0810 SB will continue to TSN; will have info for blaster.
0815 SC10 running RPG from 3 VC who are behind a barn.
0815 SA C.P. 5.
0817 Requested gunships from AAE; should be airborne in approximately fifteen minutes.

End of page 5

0820 Sb elements linked up C.P. 22.
0823 SA will be sending elements ahead to relieve road blockers.
0825 SA30 C.P. 178. SA10 and Hq from David R2.8 U0.8 SA10 clearing road block from C.P. 21 to C.P. 22.
0830 S6 chopper down.
0830 SB cleared Hoc-Mon.
0832 SC 790950
0835 Diamond Head fired on location.
0837 blaster moving to Main Gate to link up with Saber Alpha.
0845 SC20 requests artillery east of 795945
0850  SC requests artillery 795945.
0858 SA10 and 20 are going ahead to clear the road blocks.
0902 FACT will check with Saber6
0904 SB moving across SC31 front.
0905 Saber will attempt to send a medic track SC location with SA
0907 SA has linked up with Vindicator elements.
they are on the move now.
0910 Centaur40 has linked up with Centaur20 and firing.
0914 Loyal TOC reports to 2 PHs of leg soldier marching in formation at XT6623
0916 SC5 C.P. 22.

End of page 6

0919 SC5 trail C.P. 22.
0925 2 Bears tank are on way to knock out two bunkers that are putting out fire.
0927 SA10 and Hq at C.P. 8 Returning to C.P. 31
0929 centaur has contact with C31. Will attempt to drop in ammo
0930 SC5 C.P. 'D'
0932 Vindicator will shoot from XS7898.
0934 SA30 in Horse Shoe - MSR clear from C.P. 178 to C.P. 22.
0935 SC5 on arty C.P. 'E'
0938 SH101 and Medical track departing Main Gate.
0940 Tail should arrive approximately 05.
0945 SA10 back receiving BB
0945 One ship got into Charlie location and two are ready to go in.
0950 FACT on station
0952 Kind Month 9 reports VC digging in 504202
0952 Kind Month 9 reports road blown south of C.P.31 SA reports contact south of C.P.31.
0953 2 dirt mounds in route. 2 dustoffs to 94th EVAC Hosp.
0956 1 of resupply ships hit, giving ok
0959 SA30 cleared C.P.21, enroute to C.P.31
1000 Vindicator going into position.

End of page 7

1009 Kind Month 7 reports Kind month 7A and radio operator killed 2 VC between C.P. 22 and C.P. 'D'. Tropic informed.
1010 20 VC Bn in region captured enough gun powder, occupying US Embassy. 2nd Bde rendering 2nd (second) Bn to react.
1015 1 VC Co with heavy weapons 4725
1020 3 more resupply ships enroute.
1022 SB and SC linking up to provide security for evacuation.
1023 1 more resupply ship off.
1025 SA10 have artillery at location.
1030 Kind Month9 told artillery at 128, XT435248, 2 heavy weapons.
1031 1 heavy weapon Pl 4725
1035 SC5 reports 3 PC's total and 1 BB totaled.
1045 SB reports 4 more dusted off
1112 All 5 elements alert now.
1113 C15 requesting dustoff.
1115 Approximately 2 more to be dusted off.
1116 SC31. All wounded out.
1118 Sc MSR 1 convoy of TSN have taken arrival pers, both vehicles and pers.  SB has three people to be deployed west of the runway and sweep SW periodic fire 1 Pl in heavy contact.
1122 SC requested dustoff at 789952
1123 Hoc-Mon dirt trail in heavy contact.
1130 SC reports being mortared and want to know if they may be of assistance

End of page 8

1132 Go-Dau-Ha is under Red Alert
1135 SC10 location 789952
1135 Stinger has ship down on MSR
1135 SB requested all main guns be positioned.
1137 Chinook can be brought in on west end of north strip.
1140 SC going down to base about getting ground troops to organize ground sweep.
1145 SC6 is still undergoing surgery
1150 SC10 dustoff complete
1155 Mustang elements will be sweeping from north to link with SB elements
1157 2 VC Bn's still around the US Embassy and there are significant numbers of ground troops at Hoc-Mon in serious trouble; also, Tai-Khe has been under ground attack since about 0930H.
1200 SA has returned to C.P.178
1207 Stinger aircraft taken off present mission and has been diverted to another mission.
1225 S65 regained control of SC. They are now regrouping behind SB.
1230 2nd Brigade will not send any elements to Ton-Son-Nhut as there is an element (Dog Force) coming from the First Division.
1235 C31 requests dustoff of 1 man who is in shock and not wounded badly.
1240 SB35 needs dustoff. One is in contact.
1250 Report civilians are leaving Dau-Tieng.

End of page 9

1258 SC11 just picked up 3 VC POW's, all well.
1254 Centaur23 shows alarm that the ships are back up, but don't seem to have any casualties
1302 Saigon Embassy is now back in US control
1307 Centaur23 moved to Ton-Son-Nhut, some damage by .50Caliber, probably structural damage
1310 SC dustoff complete. Going to Bravo.
1317 SB10 has 10 POWs, and to Ton-Son-Nhut
1320 Corn took hit, no casualties, airstrip has no major damage.
1325 Muleskinner 504 is enroute, classified.
1326 695113. Falcon observed 1 Co. of VC.
1330 Centaur has 5 aircraft hit 4 UH1C, 2 OH23, others?.
1353 SB6 requested 90mm HE ammo.
1410 SC has a total of 11 KIA, 2 unk, N/A WIA this time.
1430 SC65 has a total of 11 KIA, 23 WIA
1432 Centaur going for dustoff at SC
1445 S OPCON to Hurricane66 (SB and SC OPCON)
1456 S6 body count 97, POW's 20.
1505 S3 to SA6. move remaining elements to GDH at 1630H.

End of page 10

1506 Kind Month9 notified that no Saber elements will be at Trang-Bang Bridge.
1535 Kind Month9 informed that no Saber elements would bo on the Hoc-Mon Bridge.
1538 SB6. 134 Body Count and 25 POW
1540 Sb65. Roster #9. Status change from WIA to KIA
1545 SB65 MED resupply needed
1550 S6 triple initial request of SB for resupply SB and SC
1600 SC6 requested dustoff for 1 WIA
1614 SH101 1 WIA, roster 93
1618 SA6 aircraft drew automatic fire 50-175 meters south of fire support base at Ton-Son-Nhut
1621 Centaur65. Centaur4 received heavy automatic weapons fire at C.P. 'E'
1628 SC POL request already into Saber4
1639 S6 - Tell Saber3 to report to Tropic3 Saber elements OPCON to Dog Face - agent of Bong-Song directed.
1640 SA10 and 30 started move now.
1648 SB65 request SA come down to SB push for resupply request.
1655 S6, SC SB night Location vic XT785930
1718 SA closed C.P. 178
1725 Fm S8. SC have 2 APC destroyed and also 1 tank was destroyed.

End of page 11

1745 Maj Hereford called Col Cowan at 1740H. Co, 725th Maint requesting pipe smoke assistance
1750 Pipe smoke airborne to Ton-Son-Nhut. Req. gunteam
1755 Centaur30 requested to know if downed Cen craft is secured by SC element - answer, affirmative.
1758 Pipe smoke of area this time.
1800 From SB65. POWs, 23 VC BC, 146
1801 To SB65. there are a large number (90 KIA). resupply will be dropped - LN90 is a KIA
1812 SB - will logger in new position; will give new logger position from Saber6
1815 SB6. to Saber 6. Orders to level textile building at 787957
1816 Pipe Smoke on final approach with gunteam
1830 S6 requests to know if SA was in contact. is being done by element OPCON LO
1832 new logger position is 810978
1833 Centaur40 is on station for new 35-40min mission
1835 Saber LP location 810978. Mission.
1836 Informed Major Deloos ADAO Centaur will be late for counter mortar.
1855 SB30 needs dustoff for 3 Killers.
1900 S-3 repts to Tropic33. Cannopt perform Dusk Patrol, - roger.

End of page 12

1906 SB reports resupply down 3 1/2 Km from their location - request to straighten it out.
1912 Dustoff airborne 1907H
1919 Cancel dustoff - dustoff already has been completed for SB30
1920 15 US KIA, 23 US WIA, 146 VC KIA BC, 23 POW's.
1922 Updated VC BC to 162.
2002 S6. SC LO 805981 where ARVNs are receiving fire at gate 005, present loc 229971
2006 Counter mortar team scrambled and DISCOM.
2026 5 airstrikes (3 ARVNs and 2 US) completed mission at textile factory.
2035 From G-2, Sgt Julian.
US Advisors Tay-Ninh reports XT443250, 1 VC Battery EVAL US B-2 ARVN B-2. XT415255 1 VC Plt B-2 ARBN and US. XT402312 and 403310 2 VC Co unevaluated.
Sent to SA, Sp4 Smith, CO.
2037 SB and SC and Vindicator at logger positions.
2045 Counter Mortar back at corral due to maintenance problems - notified ADAO
2110 From MID: Concerning and of POW and Units. 16 men - D-16 Bn 1 fm D-14, 1 fm 3d Bn, 66th Regt, 304 NVA, 1 from 6 Bn, 8th Regt. 5th Div, NVA 1 - 2nd Bn Long An Provincial Forces. 1 - 4th Bn 46th Regt, 120th NVA 1 -10th Bn, but not from C10 MR4. Movements: D16 from Tay-Ninh Prov to Me Parrats Peak to Ton-Son-Nhut. 2nd Bn Long-An from South of Duc-Hoa to Ton-Son-Nhut. 1 man D-16 personnel guess poss route of retreat along Saigon River to North, exact loc unk.

End of page 13

2120 OP XT493207 is in heavy contact with 2 Cos of VC
2130 AO Drapen and AO Tetitz cancelled
2135 Loyal loc 4/23d reports CC A Co TBB, rest at Fwd location.  Falcon all in Cu-Chi. trojan Hoc-Mon, Ton-Son-Nhut
2148 1 gunteam, 1 other available, 3d 1, 30min lift, 1 with 56, 1 just came up.
2200 TOC NE end of north runway.
2202 SA element SITREP No Change
2215 Mustang loc south and east of C.P. 'E'
2215 Inform Mustang about Killer69
2220 BC perimeter receiving fire.
2221 Mohawk reports situation mortar fire ended for OP at C.P. 31.
2222 271 Regt, 274th Regt, 275th Regt, 506th Bn, 6th Local Force, C-10 Bn, 267th Bn, 269th Bn, 2nd Tng Area Bn, possibly in region 272 and 84. F D-14 all individuals in one and moving.  The 101st and 165th as well as 273d are known to be in Saigon, but in north 2 Co. Here, 1 BN Nha-Trang, 1 BN ambushed Nha-Trang. Spread mortar fire throughout their areas.
2230 Bunker line received a few rounds and then returned fire.
2235 Go-Dau-Ha compound did receive mortars. They believe their compound is being surrounded at this time.
2240 Saber7 reports bunker line may be receiving small arms fire, but very light this time.

End of page 14

This was the last page for this Date - unknown if there are any others. 



Comments, Questions, or Suggestions send E-Mail to Webmaster
Last update on 09/11/2000 17:25